# One-loop corrections to the neutrino in the N-B-LSSM

### v1

@@ IV Numerical analysis
1. We consider the experimental constraint from the lightest CP-even Higgs mass is $125.20\pm 0.11$ GeV24s9 ; 25s33 ; 26s34 .
@@
4. The limitations for the particle masses are the following28g43 ; 29g44 ; 30g45 ; 31g46 ; 32g47 ; 33g48 . The neutralino mass is limited to more than 116 GeV, the chargino mass is limited to more than 1000 GeV and the slepton mass is greater than 600 GeV. Considering these limitations, we adopt the following parameters
@@
$~{}~{}m_{\widetilde{L}ii}^{2}=1073^{2}~{}{\rm GeV},~{}m_{\widetilde{\nu}ii}^{2}=1\times 10^{6}~{}{\rm GeV},~{}T_{Xii}=-4000~{}{\rm GeV},$
$~{}~{}T_{\nu ii}=1000~{}{\rm GeV},~{}~{}T_{eii}=1000~{}{\rm GeV}(i=1,2,3).$ (66)
we take $m^{2}_{\widetilde{E}ii}(i=1,2,3)$ as a variable, and its effect on the neutrino mass variances and three mixing angles is plotted in Fig.2.
IV.1 NO spectrum
@@
First, we study the neutrino mass spectrum at normal mass order (NO) by adjusting the model parameters. When the neutrino Yukawa coupling parameters are
$(Y_{\nu})^{11}=1.09477\times 10^{-6},~{}~{}~{}~{}~{}~{}(Y_{\nu})^{22}=1.4\times 10^{-6},$
$(Y_{\nu})^{33}=1.35146\times 10^{-6},~{}~{}~{}~{}~{}~{}(Y_{\nu})^{12}=7.6084\times 10^{-8},$
$(Y_{\nu})^{13}=4.51672\times 10^{-8},~{}~{}~{}~{}~{}~{}(Y_{\nu})^{23}=2.80238\times 10^{-7},$ (67)
@@
$\mid\Delta m^{2}_{A}\mid=0.00237154~{}{\rm eV^{2}},~{}~{}~{}\Delta m^{2}_{\odot}=0.0000799384~{}{\rm eV^{2}},$
$\sin^{2}\theta_{13}=0.00237154,~{}~{}~{}\sin^{2}\theta_{23}=0.267503,~{}~{}~{}\sin^{2}\theta_{12}=0.533371.$ (68)
Near the $3\sigma$ of neutrino detection data, we plot the diagrams.
Figure 2: The neutrino mixing angles and mass squared differences versus $m_{\widetilde{E}ii}^{2}$=ME $(i=1,2,3)$. The coloured bands in the figure represent experimental limits in the range $3\sigma$.
The parameter ME is related to the mass of the scalar lepton, which contributes to the neutrino mass matrix through the loop diagrams of the scalar lepton-chargino interactions. These corrections are crucial for understanding the small neutrino masses and mixing angles observed in experiments. Here we discuss the effect of ME, as shown in Fig.2. From the two plots in Fig.2, we can find that the neutrino mixing angle and mass variance vary weakly with respect to the ME in the range $0.5\times 10^{6}~{}{\rm GeV}^{2}-2\times 10^{6}~{}{\rm GeV^{2}}$. And in this interval, the theoretically predicted values of both mixing angle and mass variance satisfy the experimental requirements. To harmonise the other parameters, ME takes the value of $1.7\times 10^{6}~{}{\rm GeV^{2}}$(at the red line in the figure).
@@
Figure 3: The neutrino mixing angles and mass squared differences versus $({Y_{\nu}})^{33}$. The coloured bands in the figure represent experimental limits in the range $3\sigma$.
Despite the fact that neutrino Yukawa couplings coefficient are very small, they have a very significant effect on the neutrino mass and mixing angle, mainly due to their contribution at the tree level. We discuss the effect of the influence of $({Y_{\nu}})^{33}$ in Fig.3. The changes of $\Delta m_{A}^{2}$ and $\Delta m_{\odot}^{2}$ as $({Y_{\nu}})^{33}$ goes from $1.35146\times 10^{-6}$ to $4.51673\times 10^{-6}$ are demonstrated in the left panel of Fig.3. $\Delta m_{A}^{2}$ general trend of decreasing and then increasing is shown, and the two mass variances satisfy the experimental requirements at the same time around $({Y_{\nu}})^{33}=1.35146\times 10^{-6}$. The changes of $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{13}$, and $\sin^{2}\theta_{23}$ are presented in the right panels of Fig.3, and their changes are more regular in the region of $1.35144\times 10^{-6}-1.35148\times 10^{-6}$. For these three mixing angles, the appropriate region for $({Y_{\nu}})^{33}$ is around $1.35146\times 10^{-6}$.
@@
Figure 4: The neutrino mixing angles and mass squared differences versus $({Y_{\nu}})^{13}$. The coloured bands in the figure represent experimental limits in the range $3\sigma$.
In addition to the diagonal factors, the non-diagonal elements of ${Y_{\nu}}$ are also very important in our discussion. Here we discuss how $(Y_{\nu})^{13}$ affects the theoretical predictions of the neutrino mixing angle and mass-squared difference in Fig.4. Fig.4(a) shows the variation of $\Delta m_{A}^{2}$ and $\Delta m_{\odot}^{2}$. Overall, the variation $\Delta m_{\odot}^{2}$ is relatively stable when $(Y_{\nu})^{13}$ is in the range $4.6167\times 10^{-6}-4.51673\times 10^{-6}$. And the variation of $\Delta m_{A}^{2}$ is wavy. In Fig.4(a), $\sin^{2}\theta_{12}$ decreases slowly in the range of $4.6167\times 10^{-6}-4.51673\times 10^{-6}$, and $\sin^{2}\theta_{13}$ shows a general trend of decreasing and then increasing. Similarly, $\sin^{2}\theta_{23}$ shows a wavy pattern in the range of $4.6167\times 10^{-6}-4.51673\times 10^{-6}$. The theoretical values of mass variance and mixing angle around $(Y_{\nu})^{13}$ = $4.6167\times 10^{-6}$ satisfy the limits.
IV.2 IO spectrum
The neutrino mass spectrum is IO, we use the following YuKawa coupling parameters
@@ IV.2 IO spectrum
$\mid\Delta m^{2}_{A}\mid=0.00252949~{}{\rm eV^{2}},~{}~{}~{}\Delta m^{2}_{\odot}=0.0000760568~{}{\rm eV^{2}},$
$\sin^{2}\theta_{13}=0.0219436,~{}~{}~{}\sin^{2}\theta_{23}=0.311524,~{}~{}~{}\sin^{2}\theta_{12}=0.552903.$ (70)
Fig.5(a) exhibits the effect of $Y_{X11}$ on the mass variance $\Delta m_{A}^{2}$ and $\Delta m_{\odot}^{2}$ , and the right panel shows the effect of $Y_{X11}$ on the three mixing angles $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{13}$, and $\sin^{2}\theta_{23}$. In Fig.5(b), the trend of $m_{\odot}^{2}$ is shown to be relatively stable, and small changes in $Y_{X11}$ in the range $0.43-0.443$ have a significant effect on $m_{A}^{2}$. In Fig.5(b), it is shown that the value of $\sin^{2}\theta_{23}$ remains stable in the $Y_{X11}$ range $0.43-0.443$. $\sin^{2}\theta_{13}$ overall decreases slightly and then tends to stabilise. $\sin^{2}\theta_{12}$ decreases rapidly and then increases slowly in a wave-like manner. All three mixing angles near $Y_{X11}$=0.434 conform to the experimental limits.
Figure 5: The neutrino mixing angles and mass squared differences versus $Y_{X11}$. The coloured bands in the figure represent experimental limits in the range $3\sigma$.
@@
Figure 6: The neutrino mixing angles and mass squared differences versus $({Y_{\nu}})^{33}$. The coloured bands in the figure represent experimental limits in the range $3\sigma$.
@@
Figure 7: The neutrino mixing angles and mass squared differences versus $({Y_{\nu}})^{13}$. The coloured bands in the figure represent experimental limits in the range $3\sigma$.

### v2

@@ IV Numerical analysis
1. We consider that the experimental constraint from the lightest CP-even Higgs mass is $125.20\pm 0.11$ GeV24s9 ; 25s33 ; 26s34 .
@@
4. The limitations for the particle masses are the following28g43 ; 29g44 ; 30g45 ; 31g46 ; 32g47 ; 33g48 . The neutralino mass is limited to more than 116 GeV, the chargino mass is limited to more than 1100 GeV and the slepton mass is greater than 700 GeV. Considering these limitations, we adopt the following parameters
@@
$~{}~{}v_{\eta}=37.6*\cos\theta_{\eta}~{}{\rm TeV},~{}~{}v_{\overline{\eta}}=37.6*\sin\theta_{\eta}~{}{\rm TeV},~{}~{}\tan\theta_{\eta}=0.62,$
$~{}~{}m_{\widetilde{L}ii}^{2}=1073^{2}~{}{\rm GeV^{2}},~{}m_{\widetilde{\nu}ii}^{2}=1\times 10^{6}~{}{\rm GeV^{2}},~{}T_{Xii}=-4000~{}{\rm GeV},$
$~{}~{}T_{\nu ii}=1000~{}{\rm GeV},~{}~{}T_{eii}=1000~{}{\rm GeV},~{}~{}Y_{Xii}=0.43,$
$~{}~{}m_{\widetilde{E}ii}^{2}=1.7\times 10^{6}~{}{\rm GeV^{2}}~{}~{}(i=1,2,3).$ (66)
In this study, the numerical results imply that the one-loop corrections contribute approximately 10% to the total results including tree and one-loop contributions. Here we show the current experimental uncertainty and the future sensitivity in the table 2exp1 ; Jiangmen ; DaYa , where we can find the current experimental uncertainties of neutrino mass squared difference and mixing angles vary from $1.1\%$ to $4.2\%$. The corresponding future sensitivities are proposed in the approximated region $[0.2\%,~{}3.4\%]$. In the end, one can find that the one-loop corrections are about $3\sim 5$ times as the current experiment sensitivities. Therefore, it is important to consider the one loop corrections to the neutrino mass and mixing.
Table 2: Current experimental uncertainties and future sensitivitiesexp1 ; Jiangmen ; DaYa .
parameters Ordering PDG2024 $1\sigma(\%)$ Prospect(%, years)
$\frac{\Delta m_{21}^{2}}{10^{-5}{\rm eV}^{2}}$ NO, IO $7.53\pm 0.18$ 2.4 (0.3, 6)
$\frac{|\Delta m_{32}^{2}|}{10^{-3}{\rm eV}^{2}}$ NO $2.455\pm 0.028$ 1.1 (0.2, 6)
IO $2.529\pm 0.029$ 1.1
$\sin^{2}\theta_{12}$ NO, IO $0.307\pm 0.013$ 4.2 (0.5, 6)
$\frac{\sin^{2}\theta_{13}}{10^{-2}}$ NO, IO $2.19\pm 0.07$ 3.2 (2.9, -)
$\sin^{2}\theta_{23}$ NO $0.558^{+0.015}_{-0.024}$ 3.6 (0.7-3.4, 10)
IO $0.553^{+0.016}_{-0.024}$ 3.7
The main effect is due to the tree-level seesaw contribution. We want to obtain an approximate analytic understanding of the result at tree level. It would be good to explain the numerical results using analytic estimates. So, we perform approximate analytic analysis at normal order and inverse order.
IV.1 Approximate analysis of the mass matrix at tree level (NO)
At first, we discuss the approximate analytic understanding for the NO condition. Without any setting and approximation, the analytic results will be very complicated and impossible to read. Therefore, we use the following relations to simplify the discussion
$(Y_{\nu})^{11}=-1.8\times(Y_{\nu})^{22},~{}~{}(Y_{\nu})^{23}=2.6\times(Y_{\nu})^{22},~{}~{}(Y_{\nu})^{12}=(Y_{\nu})^{13}=1.5\times(Y_{\nu})^{33}.$ (67)
Then the remaining two variables of the Yukawa coupling elements $(Y_{\nu})^{ij}$ with $i,~{}j=1,~{}2,~{}3$ are $(Y_{\nu})^{22}$ and $(Y_{\nu})^{33}$. With the values for $v_{u},~{}v_{\eta},~{}Y_{X}$ in Eq.(37), when we adjust the values of the both variables $(Y_{\nu})^{22}$ and $(Y_{\nu})^{33}$, $\phi$ is a small value, stable around 0.0267. So, we simplify the formulas for neutrino mass in Eqs.(33)(35) with the assumption $\sin\phi\sim\phi\sim 0.0267$ and $\cos\phi\sim 1$,
$m_{1}^{2}=\frac{a}{3}-\frac{1}{3}p(1+\sqrt{3}*0.0267),$
$m_{2}^{2}=\frac{a}{3}-\frac{1}{3}p(1-\sqrt{3}*0.0267),$
$m_{3}^{2}=\frac{a}{3}+\frac{2}{3}p,$
$\Delta m_{\odot}^{2}=m_{\nu_{2}}^{2}-m_{\nu_{1}}^{2}={2\over\sqrt{3}}p*0.0267,$
$\Delta m_{A}^{2}=m_{\nu_{3}}^{2}-m_{\nu_{1}}^{2}=p(1+{1\over\sqrt{3}}*0.0267).$ (68)
In this condition, $a$ and $p$ are
$a=\frac{1}{2v_{\eta}^{2}Y_{X}^{2}}(-129.933\epsilon_{22}^{4}-27.04\epsilon_{22}^{3}\epsilon_{33}-162.4\epsilon_{22}^{2}\epsilon_{33}^{2}-30.6\epsilon_{22}\epsilon_{33}^{3}-50.5\epsilon_{33}^{4}),$
$p=\frac{1}{2v_{\eta}^{2}Y_{X}^{2}}\Big{(}6856.37\epsilon_{22}^{8}+9882.19\epsilon_{22}^{7}\epsilon_{33}+22477.8\epsilon_{22}^{6}\epsilon_{33}^{2}+20979.9\epsilon_{22}^{5}\epsilon_{33}^{3}$
$+22631.7\epsilon_{22}^{4}\epsilon_{33}^{4}+13669.1\epsilon_{22}^{3}\epsilon_{33}^{5}+8704.7\epsilon_{22}^{2}\epsilon_{33}^{6}+2811.15\epsilon_{22}\epsilon_{33}^{7}+1016.31\epsilon_{33}^{8}\Big{)}^{1/2},$ (69)
with $\epsilon_{ij}=\frac{v_{u}}{\sqrt{2}}(Y_{\nu})^{ij}$ and $Y_{X11}\simeq Y_{X22}\simeq Y_{X33}\simeq Y_{X}$.
As $(Y_{\nu})^{22}=1.4\times 10^{-6}$ and $(Y_{\nu})^{33}=5.0\times 10^{-7}$, the approximate numerical results are
$m_{\nu_{1}}\simeq 1.743\times 10^{-2}~{}{\rm eV},~{}~{}~{}m_{\nu_{2}}\simeq 1.997\times 10^{-2}~{}{\rm eV},~{}~{}~{}m_{\nu_{3}}\simeq 5.863\times 10^{-2}~{}{\rm eV},$
$\Delta m_{\odot}^{2}=9.513\times 10^{-5}~{}{\rm eV}^{2},~{}~{}~{}\Delta m_{A}^{2}=3.133\times 10^{-3}~{}{\rm eV}^{2}.$ (70)
IV.2 Approximate analysis of the mass matrix at tree level (IO)
In the similar way, we study the tree level neutrino mass in the IO condition with following relations
$(Y_{\nu})^{12}=-2.0\times(Y_{\nu})^{22},~{}~{}~{}~{}(Y_{\nu})^{11}=1.4\times(Y_{\nu})^{22},$
$(Y_{\nu})^{33}=-0.2\times(Y_{\nu})^{23},~{}~{}~{}~{}(Y_{\nu})^{13}=1.5\times(Y_{\nu})^{23}.$ (71)
The remaining variables are $(Y_{\nu})^{22}$ and $(Y_{\nu})^{23}$, with the simplified results of $a$ and $p$
$a=\frac{1}{2v_{\eta}^{2}Y_{X}^{2}}(-106.602\epsilon_{22}^{4}-16.04\epsilon_{22}^{2}\epsilon_{23}^{2}-1.48\epsilon_{22}\epsilon_{23}^{3}-21.6466\epsilon_{23}^{4}),$ (72)
$p=\frac{1}{2v_{\eta}^{2}Y_{X}^{2}}\Big{(}11226.8\epsilon_{22}^{8}+855.576\epsilon_{22}^{6}\epsilon_{23}^{2}-314.412\epsilon_{22}^{5}\epsilon_{23}^{3}-1578.84\epsilon_{22}^{4}\epsilon_{23}^{4}$
$\hskip 19.91684pt-63.6909\epsilon_{22}^{3}\epsilon_{23}^{5}+314.746\epsilon_{22}^{2}\epsilon_{23}^{6}+110.971\epsilon_{22}\epsilon_{23}^{7}+133.876\epsilon_{23}^{8}\Big{)}^{1/2}.$ (73)
In this condition, $\phi$ is very near $\frac{\pi}{3}$ with the remaining two variables $(Y_{\nu})^{22}$ and $(Y_{\nu})^{23}$. To deal with $\Delta m_{\odot}^{2}$ better, we adopt $\phi=\frac{\pi}{3}-\delta\phi$, and obtain the following approximate results keeping the first order of $\delta\phi$.
$m_{1}^{2}={a\over 3}-{2\over 3}p,$
$m_{2}^{2}={a\over 3}-{1\over 3}p(-1+\sqrt{3}\delta\phi),$
$m_{3}^{2}={a\over 3}+{2\over 3}p(\frac{1}{2}+\frac{\sqrt{3}}{2}\delta\phi),$
$\Delta m_{\odot}^{2}=m_{\nu_{2}}^{2}-m_{\nu_{1}}^{2}=p\frac{2\delta\phi}{\sqrt{3}},$
$\Delta m_{A}^{2}=m_{\nu_{3}}^{2}-m_{\nu_{2}}^{2}=-p(1+{1\over\sqrt{3}}\delta\phi).$ (74)
Here, we use $(Y_{\nu})^{22}=1.4\times 10^{-6}$ and $(Y_{\nu})^{23}=2.302\times 10^{-6}$, then $\delta\phi$ is around 0.024. The numerical results read as
$m_{\nu_{1}}\simeq 6.139\times 10^{-2}~{}{\rm eV},~{}~{}~{}m_{\nu_{2}}\simeq 6.147\times 10^{-2}~{}{\rm eV},~{}~{}~{}m_{\nu_{3}}\simeq 2.169\times 10^{-2}~{}{\rm eV},$
$\Delta m_{\odot}^{2}=9.154\times 10^{-5}~{}{\rm eV}^{2},~{}~{}~{}\Delta m_{A}^{2}=3.348\times 10^{-3}~{}{\rm eV}^{2}.$ (75)
IV.3 NO spectrum numerical analysis
To simplify the calculations, we have assumed that the Yukawa coupling matrix is a symmetric matrix $(Y_{\nu}=Y_{\nu}^{T})$. This assumption reduces the number of independent parameters from 9 to 6. We further assume that $Y_{\nu}$ is a real matrix. The both assumptions do not affect the calculation substantively, which simplifies the computation.
@@ IV.3 NO spectrum numerical analysis
Firstly, we study the neutrino mass spectrum at normal mass order (NO) by adjusting the model parameters. When the neutrino Yukawa coupling parameters are
$(Y_{\nu})^{11}=-2.5143\times 10^{-6},~{}~{}~{}~{}~{}~{}(Y_{\nu})^{22}=1.4\times 10^{-6},$
$(Y_{\nu})^{33}=5.1348\times 10^{-7},~{}~{}~{}~{}~{}~{}~{}~{}(Y_{\nu})^{12}=7.6533\times 10^{-7},$
$(Y_{\nu})^{13}=7.9732\times 10^{-7},~{}~{}~{}~{}~{}~{}~{}~{}(Y_{\nu})^{23}=3.6832\times 10^{-6}.$ (76)
@@
$\Delta m^{2}_{A}=2.47\times 10^{-3}~{}{\rm eV^{2}},~{}~{}~{}\Delta m^{2}_{\odot}=7.89\times 10^{-5}~{}{\rm eV^{2}},$
$m_{\nu_{1}}=1.52\times 10^{-2}~{}{\rm eV},~{}~{}m_{\nu_{2}}=1.76\times 10^{-2}~{}{\rm eV},~{}~{}m_{\nu_{3}}=5.20\times 10^{-2}~{}{\rm eV},$
$\sin^{2}\theta_{13}=0.0225,~{}~{}~{}\sin^{2}\theta_{12}=0.268,~{}~{}~{}\sin^{2}\theta_{23}=0.562.$ (77)
We scan pure loop-level parameters such as $g_{YB}$, whose influence to the one-loop corrections is approximately at the 10% level. One-loop corrections account for approximately 10% of the tree level results. Therefore, in the summary, the imapct of $g_{YB}$ is estimateed to be around 1% of the total results. Due to this relatively minor variation, it’s hard to see the change of the result in the graph. So, near the $3\sigma$ region of neutrino experiment data, we choose the other parameters as variables to plot figures in the following.
@@
Figure 2: The neutrino mass-squared differences and mixing angles versus $Y_{X11}$. The coloured bands in the figure represent experimental limits in the range of $3\sigma$.
$Y_{X}$ directly generates the right-handed neutrino mass term and indirectly modulates loop corrections to the neutrino mass matrix. Fig.2 examines the impact of $Y_{X11}$ on neutrino mass-squared differences and mixing angles. Fig.2(a) displays the evolution of $\Delta m_{A}^{2}$ and $\Delta m_{\odot}^{2}$ across $Y_{X11}\in[0.25,0.5]$. The mass-squared difference $\Delta m_{A}^{2}$ tends to change steadily after slowly decrease. In contrast, $\Delta m_{\odot}^{2}$ exhibits a clear decline trend then tends to be stable. Both $\Delta m_{\odot}^{2}$ and $\Delta m_{A}^{2}$ converge within experimental constraints (gray band) for $Y_{X11}\in[0.419,0.45]$. Fig.2(b) displays the $Y_{X11}$ effects ($0.35$ to $0.5$) on mixing angles $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{13}$, and $\sin^{2}\theta_{23}$. Both $\sin^{2}\theta_{13}$ and $\sin^{2}\theta_{23}$ demonstrate a gradual increasing trend. Conversely, $\sin^{2}\theta_{12}$ decreases clearly with increasing $Y_{X11}$. Crucially, all mixing angles simultaneously satisfy experimental limits within $Y_{X11}\in[0.425,0.442]$ (gray region). To ensure consistency with other parameters, we take $Y_{X11}=0.43$ as the benchmark value.
@@
Figure 3: The neutrino mass-squared differences and mixing angles versus $({Y_{\nu}})^{33}$. The coloured bands in the figure represent experimental limits in the range of $3\sigma$.
Despite the fact that neutrino Yukawa couplings coefficient are very small, they have a very significant effect on the neutrino mass and mixing angle, mainly due to their contribution at the tree level. Fig. 3(a) shows the evolution of the mass-squared differences ($\Delta m_{A}^{2}$ and $\Delta m_{\odot}^{2}$) across a scan of $(Y_{\nu})^{33}$ from $3\times 10^{-7}$ to $7\times 10^{-7}$. Over this domain, $\Delta m_{A}^{2}$ undergoes a slight rise. Conversely, $\Delta m_{\odot}^{2}$ shows a significant decrease. Crucially, both fall within experimentally allowed ranges (gray band) for $(Y_{\nu})^{33}$ values between $5\times 10^{-7}$ and $5.4\times 10^{-7}$. The corresponding variations in the neutrino mixing angles are plotted in Fig.3(b). $\sin^{2}\theta_{13}$ decreases with increasing $(Y_{\nu})^{33}$, though the magnitude of change is marginal. $\sin^{2}\theta_{23}$ is also decreasing but the variation is slightly larger than that of $\sin^{2}\theta_{13}$. In contrast, $\sin^{2}\theta_{12}$ experiences a consistent upward change. At the benchmark point $(Y_{\nu})^{33}=5.1348\times 10^{-7}$, all neutrino mass-squared differences and mixing angles concurrently align with established experimental limits.
Figure 4: The neutrino mass-squared differences and mixing angles versus $({Y_{\nu}})^{13}$. The coloured bands in the figure represent experimental limits in the range of $3\sigma$.
In addition to the diagonal factors, the non-diagonal elements of ${Y_{\nu}}$ are also very important in our discussion. We investigate the impact of $(Y_{\nu})^{13}$ on theoretical predictions for neutrino mass-squared differences and mixing angles in Fig.4. Fig.4(a) displays the evolution of $\Delta m_{A}^{2}$ and $\Delta m_{\odot}^{2}$ with varying $(Y_{\nu})^{13}$. The mass-squared difference $\Delta m_{A}^{2}$ exhibits a stable, near-linear rise across the parameter range. Conversely, $\Delta m_{\odot}^{2}$ demonstrates a two-phase evolution: first showing a slight decrease followed by a significant upward trend. Both quantities converge within experimental bounds (gray band) in the vicinity of $(Y_{\nu})^{13}=7.9732\times 10^{-7}$. The corresponding mixing angle variations are shown in Fig.4(b). $\sin^{2}\theta_{13}$ exhibits a gradual increase and $\sin^{2}\theta_{23}$ exhibits fairly stable over the scanned range. In contrast, $\sin^{2}\theta_{12}$ undergoes a different characteristic: decreasing initially before trending upward at around $(Y_{\nu})^{13}=6.8\times 10^{-7}$. Remarkably, at the benchmark value $(Y_{\nu})^{13}=7.9732\times 10^{-7}$, the theoretical predictions for mass-squared differences and mixing angles simultaneously satisfy the experimental constraints.
IV.4 IO spectrum numerical analysis
If the neutrino mass spectrum is IO, we use the following Yukawa coupling parameters
@@ IV.4 IO spectrum numerical analysis
$\mid\Delta m^{2}_{A}\mid=2.53\times 10^{-3}~{}{\rm eV^{2}},~{}~{}~{}\Delta m^{2}_{\odot}=7.605\times 10^{-5}~{}{\rm eV^{2}},$
$m_{\nu_{1}}=5.29\times 10^{-2}~{}{\rm eV},~{}~{}~{}m_{\nu_{2}}=5.36\times 10^{-2}~{}{\rm eV},~{}~{}~{}m_{\nu_{3}}=1.86\times 10^{-2}~{}{\rm eV},$
$\sin^{2}\theta_{13}=0.0219,~{}~{}~{}\sin^{2}\theta_{12}=0.311,~{}~{}~{}\sin^{2}\theta_{23}=0.552.$ (79)
Fig.5(a) exhibits the effect of $Y_{X11}$ on the mass variances $\Delta m_{A}^{2}$ and $\Delta m_{\odot}^{2}$, and the right panel shows the effect of $Y_{X11}$ on the three mixing angles $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{13}$, and $\sin^{2}\theta_{23}$. In Fig.5(b), the trend of $\Delta m_{\odot}^{2}$ is shown to be relatively stable, and small changes in $Y_{X11}$ in the range $0.43-0.443$ have a significant effect on $\Delta m_{A}^{2}$. In Fig.5(b), it is shown that the value of $\sin^{2}\theta_{23}$ remains stable in the $Y_{X11}$ range $0.43-0.443$. $\sin^{2}\theta_{13}$ overall decreases slightly and then tends to stabilise. $\sin^{2}\theta_{12}$ decreases rapidly and then increases slowly in a wave-like manner. All three mixing angles near $Y_{X11}$=0.434 conform to the experimental limits.
Figure 5: The neutrino mass-squared differences and mixing angles versus $Y_{X11}$. The coloured bands in the figure represent experimental limits in the range of $3\sigma$.
@@
Figure 6: The neutrino mass-squared differences and mixing angles versus $({Y_{\nu}})^{33}$. The coloured bands in the figure represent experimental limits in the range of $3\sigma$.
@@
Figure 7: The neutrino mass squared differences and mixing angles versus $({Y_{\nu}})^{13}$. The coloured bands in the figure represent experimental limits in the range of $3\sigma$.